OpenStudy (loser66):

PDE help, please Solve \(tU_x - xU_t=0\\U(x,0)= 2x^2\)

OpenStudy (loser66):

\(-\dfrac{t}{x}U_x+U_t =0\) Seek \((x(t),t)\) such that \(U(x(t),t)\) is a constant along it. characteristics equation \(\dfrac{\partial U}{\partial t}=\dfrac{\partial U}{\partial x}\dfrac{\partial x}{\partial t}+\dfrac{\partial U}{\partial t}=0\) Make a comparison to the problem, I have \(\dfrac{\partial x}{\partial t}=\dfrac{-t}{x}\) Which gives me \(\dfrac{x^2}{2}=\dfrac{-t^2}{2}+ x_0 (constant)\)

OpenStudy (loser66):

Hence \(2x_0= x^2+t^2\)